OpenStudy (anonymous):

HELP 5√2+√18 A.8 √2 B.6√18 C.14 √2 D.18 √2

OpenStudy (alexalas):

18/2=?

OpenStudy (anonymous):

9

OpenStudy (alexalas):

square root (9)=?

OpenStudy (anonymous):

81

OpenStudy (alexalas):

no that is square 9 square root 9 is 3

OpenStudy (anonymous):

i forgot

OpenStudy (alexalas):

\[\sqrt{18}=3\sqrt{2}\] so \[5\sqrt{2}+3\sqrt{2}=?\]

OpenStudy (anonymous):

18 √2

OpenStudy (alexalas):

no think of it like fractions you have to get the bottom parts to match like 1/2+2/4 you reduce or factor 2/4 to 1/2. so now it is 1/2+1/2 you add them together to get 2/2 Now with square roots it would be like \[5\sqrt{5}+4\sqrt{125}=5\sqrt{5}+4\sqrt{25*5}=5\sqrt{5}+(4+5)\sqrt{5}=5\sqrt{5}+(9)\sqrt{5}=(5+9)\sqrt{5}=14\sqrt{5}\] if that makes sense

OpenStudy (anonymous):

so then i am confused

OpenStudy (anonymous):

\[5 \sqrt{2}+\sqrt{18}=5 \sqrt{2}+3 \sqrt{2}=8 \sqrt{2} \]

OpenStudy (anonymous):

tanks

OpenStudy (anonymous):

Welcome.